OpenStudy (anonymous):

prove that [1^{2}+3^{2}+...+(2n-1)^{2}=(4n ^{3}-n)/3] for all natural number n

OpenStudy (turingtest):

step 1) see if it is true for \(n=1\)

OpenStudy (turingtest):

\[1^2={4(1)^3-(1)\over3}\]\[\LARGE\checkmark\]step 2) assume it is true for some \(n\)...

OpenStudy (turingtest):

\[\sum_{i=0}^n(2i+1)^2={4n-n\over3}\]step 3) see if this implies that it is true for \(n+1\)

OpenStudy (turingtest):

sorry that should be\[1^2+3^2+...+(2n-1)^2={4n^3-n\over3}\]the next step is then proving this implies it is true for \(n+1\)\[1^2+3^2+...+(2n-1)^2+[2(n+1)-1]^2\]\[=1^2+3^2+...+(2n-1)^2+(2n+1)^2\]\[={4n^3-n\over3}+(2n+1)^2\]you will know that you have proven it if you can show that this last line equals\[4(n+1)^3-(n+1)\over3\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

did not work

OpenStudy (turingtest):

it does work...\[\frac{4n^3-n}3+(2n+1)^2\]\[\frac{4n^3-n}3+4n^2+4n+1\]\[\frac{4n^3+12n^2+12n-n+3}3\]\[\frac{4n^3+12n^2+12n+4-n-1}3\]\[\frac{4(n^3+3n^2+3n+1)-(n+1)}3\]\[\frac{4(n+1)^3-(n+1)}3~~~~\huge\checkmark\]